find the limit at infinity of (2+x)30 x (4+x)5 / (2-x)35
1
Expert's answer
2021-09-06T14:27:14-0400
"\lim\limits_{x\to\infin}\dfrac{(2+x)^{30}\times(4+x)^5}{(2-x)^{35}}"

"=\lim\limits_{x\to\infin}\dfrac{(\dfrac{2}{x}+\dfrac{x}{x})^{30}\times(\dfrac{4}{x}+\dfrac{x}{x})^5}{(\dfrac{2}{x}-\dfrac{x}{x})^{35}}"

"=\lim\limits_{x\to\infin}\dfrac{(\dfrac{2}{x}+1)^{30}\times(\dfrac{4}{x}+1)^5}{(\dfrac{2}{x}-1)^{35}}"

"=\dfrac{(0+1)^{30}\times(0+1)^5}{(0-1)^{35}}"

"=-1"